OpenStudy (anonymous):

Simplify: 10a^2 b^3 + 30ab^2 – 20ab^3 / 5ab^2

OpenStudy (anonymous):

\[10a^2b^3 + 30ab^2 - \frac{20ab^3}{5ab^2} = 10a^2b^3 + 30ab^2 - 4b \] \[= 2b(10(ab)^2 + 15ab - 2)\] and that can't be factored as a quadratic I don't think...

jhonyy9 (jhonyy9):

(10a^2b^3 +30ab^2 -20ab^3)/5ab^2 = 5ab^2(2ab +6 -4b)/5ab^2 = =2ab +6 -4b

OpenStudy (anonymous):

yeah that looks like it, pity the question didn't make use of brackets where necessary...